### v1

@@ 4.1 Problem Instances and Training
Fleet configurations represent modern electric vehicle specifications with battery capacities of 20 and 40 kWh, fleet sizes of 4 and 8 vehicles, and ride-sharing capacities of 1, 2, and 3 passengers. Charging dynamics follow the nonlinear charging curve defined in Eq. (5), based on empirical data from [wassiliadis2021review]. Ten charging stations are placed at fixed locations in the Downtown and Civic Center districts for consistency across instances. Time windows are set to 15 minutes for both pickups and deliveries, providing moderate flexibility while maintaining realistic service quality requirements.
@@ 5.1 Case Study 1: Benchmark Performance
This case study evaluates solution quality and computational efficiency on standard benchmark instances, comparing GREAT against three exact optimization methods: column generation (CG), branch-and-price (B&P), and branch-and-cut (B&C) from [su2024branch]. Table 2 presents solution quality gaps relative to best-known solutions alongside computation times.
Table 2: Solution quality and computation time comparison on benchmark instances. Methods compared: Column Generation (CG), Branch & Price (B&P), Branch & Cut (B&C), and GREAT. Gap computed as percentage deviation from best-known solution: $\text{Gap}\%=100\times(z-z^{*})/z^{*}$. Exact methods terminated at 7,200-second limit for u5-50.
CG B&P B&C GREAT
Instance Gap% Time (s) Gap% Time (s) Gap% Time (s) Gap% Time (s)
u2-16 (16 requests) 0.92 43.6 0.00 22.2 0.00 21.0 0.00 <<1.0
u3-30 (30 requests) 0.00 570.8 0.00 570.8 0.00 438.0 0.00 <<1.0
u5-50 (50 requests) 0.05 7200.0 0.05 7200.0 7.09 7200.0 0.40 <<1.0
@@
GREAT achieves optimal or near-optimal solutions across all instances while providing substantial computational advantages. For u2-16, GREAT matches the proven optimal solutions found by B&P and B&C, whereas CG exhibits a 0.92% gap. All methods find optimal solutions for u3-30, demonstrating that the learned policy captures essential problem structure at this scale.
For u5-50, GREAT achieves a 0.40% gap compared to 0.05% for CG and B&P. While these exact methods yield marginally better solutions, B&C struggles with a 7.09% gap. GREAT thus outperforms B&C by 6.69 percentage points, indicating effective learning of routing and charging strategies. Importantly, the exact methods reached the two-hour time limit without proving optimality, whereas GREAT produces its solution in sub-second time.
@@
These results demonstrate that the neural approach scales more favorably than exact methods, which exhibit exponential growth in computation time. The consistent sub-second response time across all instance sizes confirms computational stability essential for real-time deployment. The trade-off of a small optimality gap for several orders of magnitude speedup represents a practical compromise suitable for dynamic operational environments.
@@ 5.2.1 Curriculum Learning
Computational Computation Time (s) 3.0 7200.0 0.01
@@
The RL approach demonstrates superior fleet utilization, requiring only 8.25 vehicles on average compared to 9.4 for ALNS and the full fleet of 10 for the greedy baseline. This 12% reduction in vehicles used compared to ALNS translates directly to operational cost savings. Furthermore, our method achieves a load factor of 1.94, indicating effective ridesharing with nearly two passengers transported simultaneously on average, whereas both ALNS and greedy achieve a load factor of 1, indicating single-passenger trips that underutilize vehicle capacity. The RL policy also learns efficient battery management, requiring only 0.45 charging station visits per route compared to 1.25 for ALNS and 11.8 for greedy. Despite serving all requests, RL consumes only 9.6 kWh per vehicle compared to 11.6 kWh for ALNS, demonstrating more energy-efficient routing. The greedy heuristic uses just 3.2 kWh per vehicle, but this low energy consumption reflects its poor completion rate rather than efficiency, as vehicles spend time traveling to charging stations rather than serving customers.
@@
The strong scalability of the learned policy can be attributed to several architectural and methodological factors. First, the GREAT encoder’s edge-based attention mechanism directly models the routing decision space, learning which transitions are promising based on relational features rather than absolute positions. These learned patterns—such as prioritizing requests with tight time windows or balancing vehicle workloads—generalize across problem sizes because the underlying decision logic remains consistent regardless of instance scale. Second, curriculum learning enables the policy to progressively build competence on increasingly complex instances, transferring knowledge from smaller problems rather than learning large-scale coordination from scratch. Finally, inference requires only a single forward pass through the network, yielding polynomial-time complexity compared to the exponential growth in search space faced by exact methods and the iterative evaluation required by ALNS.
@@ 5.3 Hyperparameter Sensitivity Analysis
A comprehensive hyperparameter sensitivity analysis was conducted to identify the optimal neural architecture. In total, 10 configurations were evaluated by varying three key hyperparameters: hidden dimension $d\in\{64,128,192\}$, number of attention heads $h\in\{4,8,12\}$, and number of transformer layers $L\in\{4,5,6\}$. Experiments used a 4-vehicle fleet with 20 kWh batteries and 3-passenger capacity.
@@ 5.4 Sensitivity Analysis: Reward Function Weights
Figure 4: Reward weight sensitivity analysis.
@@
These results indicate that adequate completion weight is essential for achieving full service coverage. The baseline $(1,1,1,1)$ offers balanced operation with zero lateness but no ride-sharing, while the $(1,0.1,0.1,1)$ configuration enables ride-pooling behavior through reduced wait and lateness penalties. We adopt the $(1,0.1,0.1,1)$ configuration for subsequent experiments as it achieves full completion with efficient fleet utilization through ride-sharing, and the modest lateness trade-off is acceptable for the ride-pooling scenarios considered.
@@ 5.5 Sensitivity Analysis: Battery Size, Fleet Size and Ride-Sharing Capacity
All configurations demonstrate successful convergence, with best validation profits achieved between epochs 180 and 265 and final profits ranging from 30,824 to 36,219. Four-vehicle fleets with 20 kWh batteries require the longest training (189 to 219 epochs), reflecting challenging optimization under resource constraints. Eight-vehicle fleets converge more rapidly (180 to 190 epochs), benefiting from a more forgiving optimization landscape when sufficient resources exist. The configuration with 40 kWh batteries, 4 vehicles, and three-passenger capacity requires the longest training (265 epochs), suggesting that combining moderate fleet size with high ride-sharing capacity creates a complex optimization landscape that warrants extended training budgets. Having established convergence across all configurations, we now examine how each parameter affects operational performance.
@@
Fleet size emerges as the dominant factor for service coverage. All 8-vehicle configurations achieve 100% completion rates regardless of battery size or ride-sharing capacity, while 4-vehicle fleets with 20 kWh batteries reach only 85.5 to 93.2% completion. Larger fleets deploy 5.04 to 6.04 vehicles on average (63–76% utilization), preserving reserve capacity for demand fluctuations, whereas 4-vehicle fleets operate at 90–100% utilization. This reserve capacity also enables charging flexibility: 8-vehicle fleets with 20 kWh batteries utilize 1.52 to 1.82 charge visits since some vehicles can charge while others maintain service. Per-vehicle energy consumption is lower in larger fleets (7.60 to 8.60 kWh for 20 kWh batteries) compared to 4-vehicle fleets (9.05 to 9.55 kWh), indicating better load balancing with service demand spread more evenly. Total fleet energy consumption is higher for 8-vehicle configurations (41.16 to 47.43 kWh versus 36.20 to 50.03 kWh for 4 vehicles), reflecting the additional energy required to achieve complete service coverage. Load factors remain similar across fleet sizes for equivalent ride-sharing capacities.
@@
Battery capacity has the greatest impact on resource-constrained fleets. For 4-vehicle configurations, increasing from 20 to 40 kWh improves completion rates from 85.5–93.2% to 99.8–100%, translating to profit improvements of 6–15%. Larger batteries enable better fleet utilization: 40 kWh configurations require fewer actively deployed vehicles (3.58 to 3.92 versus 4.00 for all 20 kWh 4-vehicle cases), as individual vehicles can serve more requests per deployment cycle. Energy consumption per vehicle increases substantially with larger batteries (12.76 to 13.07 kWh versus 9.05 to 9.55 kWh), reflecting the ability to perform more work without charging interruptions. Notably, 4-vehicle configurations with 20 kWh batteries do not charge (0.00 visits) despite incomplete service coverage. This reflects a learned trade-off under severe capacity constraints: charging would temporarily remove scarce vehicles from service, and the resulting loss in service capacity outweighs the potential benefit of extended range. In contrast, 4-vehicle fleets with 40 kWh batteries only charge when ride-sharing is enabled (3.01 visits for 3-passenger capacity), as the combination of larger battery and ride-sharing creates scenarios where mid-operation charging becomes beneficial. For 8-vehicle fleets, battery capacity has minimal impact on completion rates since sufficient vehicles already ensure full coverage, though charging patterns differ: 40 kWh configurations require fewer charge visits (0.59 to 1.65) compared to 20 kWh configurations (0.00 to 1.82).
@@
Ride-sharing capacity provides the most significant benefits in resource-constrained scenarios. For 4-vehicle fleets with 20 kWh batteries, completion rates improve from 85.5% (single occupancy) to 93.2% (three passengers), with corresponding profit gains of 9%. Load factors increase with capacity: 1.30–1.36 for two-passenger and 1.78–1.90 for three-passenger configurations, though these remain below theoretical maximums due to pickup and dropoff sequencing constraints. Ride-sharing also affects energy consumption patterns. Total fleet energy increases slightly with higher ride-sharing (36.20 to 38.20 kWh for 4-vehicle, 20 kWh configurations) as more requests are served, but for well-resourced configurations the opposite occurs: 8-vehicle fleets with 40 kWh batteries consume less total energy with ride-sharing enabled (41.16 kWh for 3 passengers versus 47.20 kWh for single occupancy) because fewer vehicle-kilometers are required when passengers share trips. For configurations that already achieve 100% completion, ride-sharing shows diminishing returns with profit improvements below 2%. The highest overall profit (36,219) is achieved with 40 kWh batteries, 8 vehicles, and three-passenger capacity, though this represents only 1.4% improvement over single occupancy (35,715). Wait times and lateness remain stable across all configurations regardless of ride-sharing capacity, with mean wait times ranging from 27 to 53 seconds and mean lateness below 31 seconds.
@@
The interactions between these parameters reveal important trade-offs for E-DARP fleet design. Service coverage depends primarily on having sufficient vehicle capacity, achievable through either larger fleets or larger batteries, but the operational characteristics differ substantially. Larger fleets provide redundancy and charging flexibility at the cost of higher total energy consumption and capital investment, while larger batteries enable efficient operation with fewer vehicles but create longer charging windows when replenishment is eventually needed. Ride-sharing acts as a multiplier on existing capacity: it substantially improves performance when resources are constrained but offers marginal gains when the fleet is already sufficient. Energy efficiency is maximized through the combination of adequate fleet size and ride-sharing, which together enable lower per-vehicle utilization and shared passenger trips. Charging behavior emerges as a learned strategy that balances service availability against range extension, with the policy avoiding charging when vehicle scarcity makes the service interruption too costly. These results offer practical guidance for fleet operators: compact cities with moderate demand can use smaller fleets with larger batteries to reduce costs, while larger cities benefit from more vehicles that allow some to charge while others serve requests. Ride-sharing helps minimize total energy consumption and provides the greatest benefit when resources are limited. The learned policies’ ability to adapt across these configurations shows their potential as planning tools for fleet sizing and vehicle selection.
@@ 6 Conclusion
Curriculum learning enables scaling to 250-request instances, where the policy outperforms ALNS by 9.5% in solution quality while achieving 100% completion and 720,000 times faster computation. The learned policy exhibits emergent ridesharing behavior with 3.5 times higher vehicle utilization than ALNS, demonstrating effective multi-passenger coordination.
@@
Several limitations suggest directions for future research. The framework assumes deterministic demand arrival patterns, whereas real-world operations face uncertain request arrivals, customer no-shows, and cancellations. Extending the methodology to handle dynamic demand through online learning or anticipatory policies would enhance operational reliability. Additionally, the formulation assumes a homogeneous fleet where all vehicles share identical battery capacity, passenger capacity, and energy consumption characteristics. Real-world fleets often comprise mixed vehicle types with varying ranges and capacities, requiring extensions to handle heterogeneous vehicle assignment and routing decisions.
This research demonstrates that edge-based deep reinforcement learning can effectively address complex vehicle routing problems with integrated operational decisions, achieving near-optimal performance with computational efficiency suitable for real-time deployment. The sub-second inference time enables deployment in dynamic environments where traditional optimization methods become computationally intractable. Beyond empirical performance, the edge-centric formulation offers a more structured representation of routing decisions by operating directly on graph edges rather than latent node embeddings. This added structural alignment with the underlying combinatorial problem may facilitate future theoretical analysis of learning dynamics, stability, and convergence properties of routing policies. While formal convergence guarantees for deep reinforcement learning remain an open challenge, the proposed architecture provides a promising foundation for such analysis. As cities worldwide pursue electrification and autonomous mobility services, the framework presented here provides both algorithmic advances and practical guidance for fleet deployment decisions.

### v2

@@ 4.2 Case Study 2: Large-Scale Realistic Instances
Fleet configurations represent modern electric vehicle specifications with battery capacities of 20 and 40 kWh, fleet sizes of 4 and 8 vehicles, and ride-sharing capacities of 1, 2, and 3 passengers. Charging dynamics follow the nonlinear charging curve defined in Eq. (5), based on empirical data from [wassiliadis2021review]. Ten charging stations are distributed across the city, with locations determined by K-means clustering on the transportation network nodes. Time windows are set to 15 minutes for both pickups and deliveries, with delivery windows offset from pickup windows by the direct travel time between origin and destination. This provides moderate flexibility while maintaining realistic service quality requirements.
@@ 5.1 Case Study 1: Benchmark Performance
This case study evaluates solution quality and computational efficiency on standard benchmark instances, comparing GREAT against four exact optimization methods: column generation (CG),column generation with cutting planes (CG-CP), branch-and-price (B&P), and branch-and-cut (B&C) from [su2024branch]. Table 2 presents solution quality gaps relative to best-known solutions alongside computation times.
Table 2: Solution quality and computation time comparison on benchmark instances. Methods compared: Column Generation (CG), Column Generation with cutting planes (CG-CP), Branch & Price (B&P), Branch & Cut (B&C), and GREAT. Gap computed as percentage deviation from best-known solution: $\text{Gap}\%=100\times(z-z^{*})/z^{*}$. Exact methods terminated at 7,200-second limit for u5-50.
CG CG-CP B&P B&C GREAT
Instance Gap% Time (s) Gap% Time (s) Gap% Time (s) Gap% Time (s) Gap% Time (s)
u2-16 (16 requests) 0.92 43.6 0.00 72.1 0.00 22.2 0.00 21.0 0.00 <<1.0
u3-30 (30 requests) 0.00 570.8 0.00 570.8 0.00 570.8 0.00 438.0 0.00 <<1.0
u5-50 (50 requests) 0.05 7200.0 0.05 7200 0.05 7200.0 7.09 7200.0 0.40 <<1.0
@@
GREAT achieves optimal or near-optimal solutions across all instances while providing substantial computational advantages. For u2-16, GREAT matches the proven optimal solutions found by CG-CP, B&P, and B&C, whereas CG exhibits a 0.92% gap. All methods find optimal solutions for u3-30, demonstrating that the learned policy captures essential problem structure at this scale.
For u5-50, GREAT achieves a 0.40% gap compared to 0.05% for CG, CG-CP, and B&P. While these exact methods yield marginally better solutions, B&C struggles with a 7.09% gap. GREAT thus outperforms B&C by 6.69 percentage points, indicating effective learning of routing and charging strategies. Importantly, the exact methods reached the two-hour time limit without proving optimality, whereas GREAT produces its solution in sub-second time.
@@
These results demonstrate that the learned policy scales more favorably than exact methods, which exhibit exponential growth in computation time. The consistent sub-second response time across all instance sizes confirms computational stability essential for real-time deployment. The trade-off of a small optimality gap for several orders of magnitude speedup represents a practical compromise suitable for dynamic operational environments.
@@ 5.2 Case Study 2: Scalability Analysis
Computational Computation Time (s) 3.0 7200.0 9.3
@@
The RL approach demonstrates superior fleet utilization, requiring only 8.25 vehicles on average compared to 9.4 for ALNS and the full fleet of 10 for the greedy baseline. This 12% reduction in vehicles used compared to ALNS translates directly to operational cost savings. Furthermore, our method achieves a load factor of 1.94, indicating effective ridesharing with nearly two passengers transported simultaneously on average, whereas both ALNS and greedy achieve a load factor of 1, indicating single-passenger trips that underutilize vehicle capacity. The RL policy also learns efficient battery management, requiring only 0.45 charging station visits per route compared to 1.25 for ALNS and 11.8 for greedy. Despite serving all requests, RL consumes only 9.6 kWh per vehicle compared to 11.6 kWh for ALNS, demonstrating more energy-efficient routing. The greedy heuristic uses just 3.2 kWh per vehicle, but this low energy consumption reflects its poor completion rate rather than efficiency.
@@
The strong scalability of the learned policy can be attributed to several architectural and methodological factors. First, the GREAT encoder’s edge-based attention mechanism directly models the routing decision space, learning which transitions are promising based on relational features rather than absolute positions. These learned patterns, such as prioritizing requests with tight time windows or balancing vehicle workloads, generalize across problem sizes because the underlying decision logic remains consistent regardless of instance scale. Second, curriculum learning enables the policy to progressively build competence on increasingly complex instances, transferring knowledge from smaller problems rather than learning large-scale coordination from scratch. Finally, inference requires only a single forward pass through the network, yielding polynomial-time complexity compared to the exponential growth in search space faced by exact methods and the iterative evaluation required by ALNS.
@@ 5.3 Hyperparameter Sensitivity Analysis
A comprehensive hyperparameter sensitivity analysis was conducted to identify the optimal neural architecture. In total, 10 configurations were evaluated by varying three key hyperparameters: hidden dimension $d\in\{64,128,192\}$, number of attention heads $h\in\{4,8,12\}$, and number of transformer layers $L\in\{4,5,6\}$. Experiments used an 8-vehicle fleet with 20 kWh batteries and 3-passenger capacity.
@@ 5.4 Sensitivity Analysis: Reward Function Weights
Figure 4: Training dynamics across reward weight configurations $(w_{e},w_{w},w_{l},w_{c})$ for 100 requests with 20 kWh batteries and 8 vehicles. Metrics shown: (a) completion rate, (b)nvehicles deployed, (c)energy consumption per vehicle, (d) customer wait time, (e) mean state of charge, and (f) vehicle load factor. High operational cost weights $(10,10,10,1)$ lead to conservative behavior and incomplete service, while reduced service penalties $(1,0.1,0.1,1)$ enable ride-sharing (load factor >1.8>1.8) at the cost of moderate lateness.
@@
These results indicate that adequate completion weight is essential for achieving full service coverage. The baseline $(1,1,1,1)$ offers balanced operation with zero lateness but no ride-sharing, while the $(1,0.1,0.1,1)$ configuration enables ride-sharing behavior through reduced wait and lateness penalties. We adopt the $(1,0.1,0.1,1)$ configuration for subsequent experiments as it achieves full completion with efficient fleet utilization through ride-sharing, and the modest lateness trade-off is acceptable for the ride-sharing scenarios considered.
@@ 5.5 Sensitivity Analysis: Battery Size, Fleet Size and Ride-Sharing Capacity
All configurations demonstrate successful convergence, with the best validation profits achieved between epochs 180 and 265 and final profits ranging from 30,824 to 36,219. Four-vehicle fleets with 20 kWh batteries require the longest training (189–219 epochs), reflecting challenging optimization under resource constraints. Eight-vehicle fleets converge more rapidly (180–190 epochs), benefiting from a more forgiving optimization landscape when sufficient resources exist. The configuration with 40 kWh batteries, 4 vehicles, and three-passenger capacity requires the longest training (265 epochs), suggesting that combining moderate fleet size with high ride-sharing capacity creates a complex optimization landscape that warrants extended training budgets. Having established convergence across all configurations, we now examine how each parameter affects operational performance.
@@
Fleet size emerges as the dominant factor for service coverage. All 8-vehicle configurations achieve 100% completion rates regardless of battery size or ride-sharing capacity, whereas 4-vehicle fleets with 20 kWh batteries reach only 85.5–93.2% completion. Larger fleets deploy 5.04–6.04 vehicles on average (63–76% utilization), preserving reserve capacity for demand fluctuations, whereas 4-vehicle fleets operate at 90–100% utilization. This reserve capacity also enables charging flexibility: 8-vehicle fleets with 20 kWh batteries perform 1.52–1.82 charge visits, as some vehicles can charge while others maintain service. Per-vehicle energy consumption is lower in larger fleets (7.60–8.60 kWh for 20 kWh batteries) than in 4-vehicle fleets (9.05–9.55 kWh), indicating better load balancing when service demand is spread more evenly. Total fleet energy consumption is higher for 8-vehicle configurations (41.16–47.43 kWh versus 36.20–50.03 kWh for 4 vehicles), reflecting the additional energy required to achieve complete service coverage. Load factors remain similar across fleet sizes for equivalent ride-sharing capacities.
@@
Battery capacity has the greatest impact on resource-constrained fleets. For 4-vehicle configurations, increasing capacity from 20 to 40 kWh improves completion rates from 85.5–93.2% to 99.8–100%, translating into profit improvements of 6–15%. Larger batteries enable better fleet utilization: 40 kWh configurations require fewer actively deployed vehicles (3.58–3.92, versus 4.00 for all 20 kWh 4-vehicle cases), as individual vehicles can serve more requests per deployment cycle. Energy consumption per vehicle increases substantially with larger batteries (12.76–13.07 kWh versus 9.05–9.55 kWh), reflecting the ability to perform more work without charging interruptions. Notably, 4-vehicle configurations with 20 kWh batteries do not charge (0.00 visits) despite incomplete service coverage. This reflects a learned trade-off under severe capacity constraints: charging would temporarily remove scarce vehicles from service, and the resulting loss in service capacity outweighs the potential benefit of extended range. In contrast, 4-vehicle fleets with 40 kWh batteries charge only when ride-sharing is enabled (3.01 visits for three-passenger capacity), as the combination of larger batteries and ride-sharing creates scenarios in which mid-operation charging becomes beneficial. For 8-vehicle fleets, battery capacity has minimal impact on completion rates since sufficient vehicles already ensure full coverage, though charging patterns differ: 40 kWh configurations require fewer charge visits (0.59–1.65) than 20 kWh configurations (0.00–1.82).
@@
Ride-sharing capacity provides the most substantial benefits in resource-constrained scenarios. For 4-vehicle fleets with 20 kWh batteries, completion rates improve from 85.5% (single occupancy) to 93.2% (three passengers), with corresponding profit gains of about 9%. Load factors increase with capacity (1.30–1.36 for two-passenger and 1.78–1.90 for three-passenger configurations), though these remain below theoretical maxima because of pickup and drop-off sequencing constraints. Ride-sharing also affects energy consumption patterns. Total fleet energy increases slightly with higher ride-sharing (36.20–38.20 kWh for 4-vehicle, 20 kWh configurations) as more requests are served, but for well-resourced configurations the opposite occurs: 8-vehicle fleets with 40 kWh batteries consume less total energy with ride-sharing enabled (41.16 kWh for three passengers versus 47.20 kWh for single occupancy) because fewer vehicle-kilometers are required when passengers share trips. For configurations that already achieve 100% completion, ride-sharing shows diminishing returns, with profit improvements below 2%. The highest overall profit (36,219) is achieved with 40 kWh batteries, 8 vehicles, and three-passenger capacity, but this represents only a 1.4% improvement over single occupancy (35,715). Wait times and lateness remain stable across all configurations, with mean wait times ranging from 27 to 53 seconds and mean lateness below 31 seconds.
@@
The interactions between these parameters reveal important trade-offs for E-DARP fleet design. Service coverage depends primarily on having sufficient vehicle capacity, achievable through either larger fleets or larger batteries, but the operational characteristics differ substantially. Larger fleets provide redundancy and charging flexibility at the cost of higher total energy consumption and capital investment, whereas larger batteries enable efficient operation with fewer vehicles but create longer charging windows when replenishment is eventually needed. Ride-sharing acts as a multiplier on existing capacity: it substantially improves performance when resources are constrained but offers marginal gains when the fleet is already sufficient. Energy efficiency is maximized through the combination of adequate fleet size and ride-sharing, which together enable lower per-vehicle utilization and shared passenger trips. Charging behavior emerges as a learned strategy that balances service availability against range extension, with the policy avoiding charging when vehicle scarcity makes service interruptions too costly. These results offer practical guidance for fleet operators: compact cities with moderate demand can use smaller fleets with larger batteries to reduce costs, whereas larger cities benefit from more vehicles that allow some to charge while others serve requests. Ride-sharing helps minimize total energy consumption and provides the greatest benefit when resources are limited. The learned policies’ ability to adapt across these configurations highlights their potential as planning tools for fleet sizing and vehicle selection.
@@ 6 Conclusion
Curriculum learning enables scaling to 250-request instances, where the policy outperforms ALNS by 9.5% in solution quality while achieving 100% completion and 774 times faster computation. The learned policy exhibits emergent ridesharing behavior with 3.5 times higher vehicle utilization than ALNS, demonstrating effective multi-passenger coordination.
@@
Several limitations suggest directions for future research. The framework assumes deterministic demand arrival patterns, whereas real-world operations face uncertain request arrivals, customer no-shows, and cancellations. Extending the methodology to handle dynamic demand through online learning or anticipatory policies would enhance operational reliability. Additionally, the formulation assumes a homogeneous fleet where all vehicles share identical battery capacity, passenger capacity, and energy consumption characteristics. Real-world fleets often comprise mixed vehicle types with varying ranges and capacities, requiring extensions to handle heterogeneous vehicle assignment and routing decisions. Furthermore, the current framework discretizes charging decisions by fixing charging duration at each station visit, whereas exact formulations optimize continuous charging times. This discretization may contribute to the small optimality gaps observed on benchmark instances (0.4% for 50-request problems) and could limit performance in scenarios requiring precise energy management. Future work could explore hybrid approaches that combine learned routing policies with optimization-based charging decisions, or action spaces that include multiple discrete charging duration options.
This research demonstrates that edge-based deep reinforcement learning can effectively address complex vehicle routing problems with integrated operational decisions, achieving near-optimal performance with computational efficiency suitable for real-time deployment. The inference time of approximately 10 seconds enables practical deployment in operational environments where traditional optimization methods requiring hours become computationally intractable. Beyond empirical performance, the edge-centric formulation offers a more structured representation of routing decisions by operating directly on graph edges rather than latent node embeddings. This added structural alignment with the underlying combinatorial problem may facilitate future theoretical analysis of learning dynamics, stability, and convergence properties of routing policies. While formal convergence guarantees for deep reinforcement learning remain an open challenge, the proposed architecture provides a promising foundation for such analysis. As cities worldwide pursue electrification and autonomous mobility services, the framework presented here provides both algorithmic advances and practical guidance for fleet deployment decisions.